OpenStudy (anonymous):

The equation of the line of best fit of a scatter plot is y = −5x − 2. What is the slope of the equation? −5 −2 2 5

OpenStudy (teddyiswatshecallsme):

Do you know the equation of slope intercept form?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

Slope intercept form is in the general form of y = mx+b If you had an equation like y = 3x+7 the slope would be m = 3 since the 'm' and '3' match up. They are both in front of the x.

jimthompson5910 (jim_thompson5910):

Other examples y = 2x+9 has slope of m = 2 y = -82x-6 has slope of m = -82 y = (2/5)x + 3 has slope of m = 2/5 y = 6.7x - 10 has slope of m = 6.7

OpenStudy (anonymous):

@jim_thompson5910 so what is the correct answer?

jimthompson5910 (jim_thompson5910):

did my examples help at all?

OpenStudy (anonymous):

@jim_thompson5910 not at all thats why i asked fir the answer so i kan see how u got it

jimthompson5910 (jim_thompson5910):

what number is in front of the x in your equation

OpenStudy (anonymous):

-5 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

thats the slope

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):